Becky is having her birthday party at a roller skating ring. The ring charges a fee of $50 plus $8 per person. If Becky invites 15 friends to her birthday party, how much will she pay for the party?


1
Expert's answer
2021-05-19T17:47:31-0400

Including becky total no of persons = 15+1=16


Hence cost = 8 × 16 = 128


Total cost = 50 + 128 = $178